# Sparse-to-Sparse Training of Diffusion Models

### v1

@@ Appendix E Experiments using Various Diffusion Timesteps
Appendix F Prune and Regrowth Rate Experiments
@@
All in all, these findings suggest that a prune and regrowth ratio of 0.5 is too aggressive, and that a more conservative choice of 0.05 is more appropriate for DMs. Previous work has mentioned that DST methods are consistently superior to SST as long as there is appropriate parameter exploration [Liu et al., 2021c], an observations that aligns with our findings.
Appendix G Examples of Generated Samples
Figures 8 to 13 showcase examples of samples generated by Latent Diffusion and ChiroDiff across the evaluated datasets. Examples are unconditionally sampled from the Dense and the top-performing sparse model in each case.

### v2

@@ Appendix G Experiments using Various Diffusion Timesteps
Appendix H Prune and Regrowth Rate Experiments
@@
All in all, these findings suggest that a prune and regrowth ratio of 0.5 is too aggressive, and that a more conservative choice of 0.05 is more appropriate for DMs. Previous work has mentioned that DST methods are consistently superior to SST as long as there is appropriate parameter exploration (Liu et al., 2021c), an observations that aligns with our findings.
Appendix I Additional evaluation metrics
In this section, we present the KID results corresponding to previously reported experiments, in Figure 2, Figure 3, Figure 9 and Figure 10.
Figure 11: KID comparisons between Dense, and Static-DM, MagRan-DM and RigL-DM with various sparsity levels, for LatentDiffusion, with prune and regrowth rate $p=0.5$.
Figure 12: KID comparisons between Dense, and Static-DM, MagRan-DM and RigL-DM with various sparsity levels, for LatentDiffusion, with prune and regrowth rate $p=0.05$.
Figure 13: KID comparisons between Dense, and Static-DM, MagRan-DM and RigL-DM with various sparsity levels, for ChiroDiff, with prune and regrowth rate $p=0.5$.
Figure 14: KID comparisons between Dense, and Static-DM, MagRan-DM and RigL-DM with various sparsity levels, for ChiroDiff, with prune and regrowth rate $p=0.05$.
Appendix J Examples of Generated Samples
Figures 15 to 20 showcase examples of samples generated by Latent Diffusion and ChiroDiff across the evaluated datasets. Examples are unconditionally sampled from the Dense and the top-performing sparse model in each case. Figures 21 and 22 present examples of class-conditional generated samples on ImageNet and QuickDraw.
@@ Appendix I Additional evaluation metrics
(a) Dense
(b) MagRan-DM, $S=0.50$
Figure 21: Samples from Latent Diffusion trained on class-conditional ImageNet. The top row presents samples generated by Dense models, whereas the bottom row presents samples generated by the top-performing sparse model.
(a) Dense
(b) MagRan-DM, $S=0.50$
Figure 22: Samples from ChiroDiff trained on class-conditional QuickDraw. The top row presents samples generated by Dense models, whereas the bottom row presents samples generated by the top-performing sparse model.